Consider two 100šœ‡š¹ capacitors. One is charged to10V, disconnected from the power supply,Ā and then connected across the other. Calculate the energy stored by the combinationĀ 


1
Expert's answer
2021-03-29T09:02:40-0400

"W=\frac{C_0 U_0^2}{2},""C_0=C_1+C_2=2C,"

"U_0=U_1=U_2=U,"

"W=CU^2,"

"W=100\cdot 10^{-6}\cdot 10^2=0.01~J."